OpenStudy (anonymous):

What is the product of five square root five times six square root four? Simplify your answer. A.) 30 square root nine B.) sixty square root five C.) sixty square root twenty-four D.) one hundred fifty square root four

OpenStudy (parker.goodbar):

5×6=? 5√×4√=?

OpenStudy (parker.goodbar):

Multiply 5 and 4.

OpenStudy (parker.goodbar):

than 5 and 6

OpenStudy (parker.goodbar):

\[30\sqrt{20}\]

OpenStudy (parker.goodbar):

than break down the 20

OpenStudy (parker.goodbar):

find gcf which is 4 and 5

OpenStudy (parker.goodbar):

so the answer is \[60\sqrt{5}\]

OpenStudy (parker.goodbar):

or B

OpenStudy (parker.goodbar):

are you still there @Nefxrious

OpenStudy (anonymous):

@parker.goodbar yes thank you! this helped me a lot and now I understand it!

OpenStudy (parker.goodbar):

anymore questions

OpenStudy (anonymous):

@parker.goodbar yes, what is the exact value for the expression the square root of 52. − the square root of 13. + the square root of 117.? Simplify if possible. i will add the answers in the next reply

OpenStudy (anonymous):

A.) four square root thirteen B.) square root thirteen C.) two square root thirty-nine D.) eight square rot thirty-nine

OpenStudy (parker.goodbar):

\[\sqrt{52}-\sqrt{13}+\sqrt{117}\]

OpenStudy (parker.goodbar):

\[2\sqrt{13}-\sqrt{13}+3\sqrt{13}\]

OpenStudy (parker.goodbar):

it is A

OpenStudy (parker.goodbar):

\[4\sqrt{13}\]

OpenStudy (parker.goodbar):

because \[2\sqrt{13}-\sqrt{13}=1\sqrt{13} or \sqrt{13}then add 3\sqrt{13}= 4\sqrt{13}\]

OpenStudy (anonymous):

@parker.goodbar how did you get 52 and 117 square rooted?

OpenStudy (parker.goodbar):

you break it down using gcf like 52 the greatest common factor that can also be rooted(wich is what you look for) is 4 wich has a square root of 2 and 52 divided by 4 equals 13 so it equals \[\sqrt{4*13} or 2\sqrt{13}\]

OpenStudy (parker.goodbar):

same with 117 except it is 9 which has a square root of 3 and divides into 117 to make 13 so the equation would be\[3\sqrt{13}\]

OpenStudy (anonymous):

@parker.goodbar thank makes a lot of sense! thank you so much for this you are a better teacher than my actual one! that is my last question thanks again!

OpenStudy (parker.goodbar):

any time i under stand my teacher sucks too have a good day or night